OpenStudy (anonymous):

Lizzie had many cookies, after eating one she gave half of the remainder to her sister. After eating another cookie, she gave half of what was left to her brother. Lizzie now only had five cookies left, how many cookies did she start with? A 12 B 17 C 21 D 23 E 43

OpenStudy (anonymous):

It should be 23. If you go the other way around, it would be 5 times 2 = 10 + 1 = 11 times 2 = 22 + 1 = 23

OpenStudy (tkhunny):

C = Original Number of Cookies. There, you're about half done. Now, translate the rest of the problem statement in terms of C.

OpenStudy (whpalmer4):

\[\frac{\frac{C-1}2 - 1}{2}=5\]

OpenStudy (anonymous):

23 is the ans if u go from end to start of the Q then 5x2=10 [add brother's half] 10+1=11 [add the one she ate] 11x2=22 [add sister's half] 22+1=23 [add the one she ate]